Two thin infinite parallel sheets (non-conducting) have uniform surface densities of charge +q and -q. Electric field in the space between the two sheets is ....
Plz mention the aanswer in terms of q/epsilon where Epsilon is the electrical permitivity.....Plz provide the answer with solution.....

Grade:12th pass

1 Answers

Ash
5 years ago
Electric field due to negavively charged sheet is
 \frac{\sigma }{2\varepsilon _{0}} towards it.
Electric field due to postively charged sheet is
 \frac{\sigma }{2\varepsilon _{0}} away from it.
 Here sigma denotes surface charge density of the sheet.
So,the net electric field is sum of those mentioned above that is
E_{net}=\frac{\sigma }{\varepsilon _{0}} towards the negatively charged sheet.